OpenStudy (anonymous):

Which property is illustrated by the statement (2+3.4)+6=2+(3.4+6)

OpenStudy (compassionate):

Hi, Handyany This property states that, no mater what form you perform the operations, you will always get the same answers. (i.g., 2 + 3 is the same thing as 3 + 2.) This property is also know as the Addition Property of Equality. You, however, have multiplication. The same rule applies (i.g., 2 x 3 is the same thing as 3 x 2.) This is also know as the Communicative Property When two numbers are multiplied together, the product is the same regardless of the order of the multiplicands. For example 4 * 2 = 2 * 4. In your example, we do not have two numbers being multiplied, however, we have three, and some grouping. When you have this, you use something called the Associate Property. Associative Property: When three or more numbers are multiplied, the product is the same regardless of the grouping of the factors. For example (2 * 3) * 4 = 2 * (3 * 4) Regards, OpenStudy Ambassador, OpenStudy Team

OpenStudy (anonymous):

Wow thank you! that's by far the best help I've ever gotten.